A consumer spends all of her income​(Y) on two goods Z and B.The price of good B ​(PB​) is $6. The Marginal Rate of Transformation MRT is equal to −2. That is 2 units of good B can be traded for 1 unit of good Z.This consumer is able to buy 18 units of good Z and 0 units of good B with​ his/her income. What is this​ consumer's level of​ income?

The​ consumer's income is ​$ ( )​(round your answer to the nearest ​penny).
Answer #1

At the optimal point,

MRS = Ratio of prices

-2 = Pz/Pb

-2 = Pz/6

Pz = $12

Next, consumer budget constraint is:

M = Pz.Z + Pb.b

M = 12(18) + 0

M = 216

Thus consumer income is $216